Consider the ANOVA table that follows.

Analysis of Variance
Source DF SS MS F
Regression 5 3710.00 742.00 12.89
Residual Error 46 2647.38 57.55
Total 51 6357.38

a-1. Determine the standard error estimate. (Round your answer to 2 decimal places.)

a-2. About 95% of the residuals will be between what two values? (Round your answers to 2 decimal places.)

b-1. Determine the coefficient of multiple determination. (Round your answer to 3 decimal places.)

b-2. Determine the percentage variation for the independent variables. (Round your answer to 1 decimal place.) Determine the coefficient of multiple determination, adjusted for the degrees of freedom. (Round your answer to 3 decimal places.)
Answer #1

a-1) standard error estimate =sqrt(57.55)=7.59

a-2) since 95% of  residuals will be between 2 standard error from mean.

95% of the residuals will be between -15.18 and 15.18

b-1)

coefficient of multiple determination =SSR/SST =3710/6357.38= 0.584

b-2)

percentage variation for the independent variables =58.4 %

coefficient of multiple determination, adjusted for the degrees of freedom R2adj=1-(1-R2)*(n-1)/(n-k-1)= 0.538